# Parametric Channel Estimation and Design for Active-RIS-Assisted Communications

### v1

@@ I Introduction
Next-generation (6G) wireless networks require innovations that reconfigure the propagation environment to improve end-to-end performance. Reconfigurable intelligent surfaces (RISs) have emerged as a promising means of enhancing wireless channels via programmable metasurfaces [1]. A conventional passive RIS comprises many low-cost, passive scatterers that reflect incident signals with adjustable phase shifts, thereby shaping the channel without additional transmit power [2]. However, passive RISs intrinsically suffer from a double-fading (multiplicative path-loss) effect: the end-to-end gain factors as the product of the transmitter-to-RIS and RIS-to-receiver gains, so the reflected signal experiences attenuation twice [3].
To overcome this limitation, researchers have proposed active RIS (ARIS) architectures in which each element integrates active components (e.g., low-power amplifiers or transceiver chips) that amplify rather than merely reflect the impinging wave [3]. By providing extra power at the surface, an active RIS can compensate for double-path loss, effectively converting multiplicative attenuation into an additive gain term in the link budget. Recent works have shown that active RIS can outperform their passive counterparts under comparable settings [1, 4]. In this sense, the active RIS concept can be viewed as a natural evolution of the traditional RIS, using a small amount of surface power to counter the effects of double fading and boost overall network throughput [1].
Even with these advantages, obtaining accurate channel state information (CSI) in RIS-aided systems is still a difficult problem. In typical RIS deployments, the base station (BS) does not observe the constituent links directly; instead, it measures a cascaded channel—the product of the user-to-RIS and RIS-to-BS channels—modulated by the RIS elements’ responses [5].
Moreover, the classical far-field assumption (planar wavefronts) may not hold for large-aperture RISs that operate in the near field, where spherical wavefront propagation must be modeled. In this regime, the channel depends not only on angular direction but also on the distance to each RIS element, introducing additional parameters [6]. Near-field channels offer higher spatial degrees of freedom. However, they substantially increase the complexity of channel modeling and estimation. Consequently, far-field (angle-only) parameterizations become inaccurate in the near field or require an excessive number of pilots  [6].
Mylonopoulos et al. [7] derived a maximum-likelihood estimator (MLE) for user position and orientation in a mmWave network with an active RIS, treating the line-of-sight (LoS) paths via the RIS as geometric signals and using maximum likelihood to infer location with high accuracy. Although their objective—localization—differs from ours, the result reinforces that MLE on parametric LoS channel models (angles and distances) is effective, and that active RIS hardware can further enhance estimation. Chen et al. [8] investigated channel estimation and training design for active RIS-aided systems in which each RIS element can amplify the reflected signal. A Cramér–Rao lower bound (CRLB) analysis shows that active amplification can reduce pilot overhead relative to passive RISs, albeit at the expense of increased noise. This early contribution proposed joint training and least-squares (LS) estimation that exploits amplification to improve estimation performance and provides insight into reflection-coefficient design. Dynamic channel tracking has recently been investigated in the context of active RISs. Haider et al. [9] proposed a semi-passive RIS architecture to handle time-varying channels (e.g., due to user mobility) while maintaining low pilot overhead. In their design, only a small subset of RIS elements are active (equipped with RF chains) and intermittently sense the channel. A long short-term memory (LSTM) network is then employed for temporal prediction; however, it requires a training phase and performs reliably only when user mobility patterns resemble those represented in the training data.
Another emerging direction employs machine learning to design efficient training protocols for active RISs. Zhang and Yu [10] developed an adaptive beamforming codebook for active RIS-aided localization; unlike conventional beam training, their approach learns codebook construction and beam selection. However, generalization may degrade when the deployment environment or hardware differs from the training data, necessitating retraining. Moreover, executing an LSTM-based selection loop in real time increases controller complexity, as it must perform on-the-fly neural inference.
In this letter, we exploit the signal-amplification capability of active reconfigurable intelligent surfaces (ARISs) to tackle the challenging problem of channel estimation. We develop a parametric maximum likelihood estimation (MLE) framework and derive closed-form expressions for all channel parameters under both near-field and far-field conditions.
Compared with passive RISs, channel estimation for active RISs is complicated by two key factors: (i) amplified thermal and phase noise, which obscures parameter identification, and (ii) practical implementation requirements, as each active element demands precise calibration of both phase and amplitude. We address these challenges with a methodology tailored to the active RIS architecture that jointly estimates the amplitude and phase components of the channel response. The proposed framework delivers closed-form localization parameters—angles for far-field users and angles plus distance for near-field users—and includes a beamforming design specifically adapted to active RIS hardware.
A central feature of the framework is an adaptive pilot-selection mechanism that dynamically allocates pilots to the most informative directions as estimation proceeds, improving accuracy without increasing pilot overhead. To the best of our knowledge, this is the first comprehensive parametric estimation solution explicitly designed for the amplification and noise characteristics of active RISs, and it operates seamlessly across both near-field and far-field regimes.
@@ II System and Channel Model
We consider an uplink system comprising a single-antenna base station (BS) and a single-antenna user equipment (UE) assisted by an active reconfigurable intelligent surface (ARIS). No direct BS-to-UE link is assumed. The active RIS comprises $N$ elements with independently tunable amplitude and phase. We model the BS-to-ARIS channel, $\bm{h}\in\mathbb{C}^{N}$, as a deterministic Line-of-Sight (LoS) link. Given that the BS and active RIS are fixed in location, this channel is considered quasi-stationary and non-fading over the signaling interval. The ARIS-to-UE channel is $\bm{g}\triangleq[g_{1},\ldots,g_{N}]^{\mathrm{T}}\in\mathbb{C}^{N}$, which is time varying due to user mobility. To maintain a high signal-to-noise ratio (SNR), the BS periodically estimates the channel and updates the active RIS coefficients (phase shifts and gains) accordingly.
Let the UE transmit the symbol $x\sim\mathcal{CN}(0,P_{d})$. The received signal at the BS is
@@
where $\bm{\Phi}\triangleq[\phi_{1},\ldots,\phi_{N}]^{\mathrm{T}}$ , it collects the complex active RIS coefficients (per-element gain and phase). Here $\mathbf{D}_{h}=\operatorname{diag}(h_{1},\ldots,h_{N})$, and $\Phi_{n}=p_{n}e^{\mathrm{j}\theta_{n}}$ is the active RIS configuration, where $\theta_{n}$ is the phase shift and $p_{n}$ is the amplification factor of the $n$th element, respectively.
@@
The spectral efficiency (SE) can be expressed as
@@
The line-of-sight (LoS) channel vector from UE-to-ARIS is modeled as
$\bm{g}\;=\;\sqrt{\beta}\,e^{\mathrm{j}\omega}\,\bm{a}(\psi)$ (3)
where $\beta\geq 0$ and $\omega\in[0,2\pi)$ denote, respectively, the channel gain and the phase referenced to the first active RIS element, and $\bm{{a}(\psi)}$ is the array-response vector parameterized by $\bm{\psi}$. This model applies to both near- and far-field propagation: in the far field, $\bm{\psi}=(\varphi,\phi)$ comprises the azimuth $\varphi$ and elevation $\phi$; in the near field, $\bm{\psi}=(\varphi,\phi,r)$ additionally includes the distance $r$ from the UE to the first active RIS element.
For the near field, the array-response vector can be written as [11]
$\mathbf{a}(\varphi,\phi,r)=\begin{bmatrix}1,\,e^{\mathrm{j}\frac{2\pi}{\lambda}s_{2}},\,e^{\mathrm{j}\frac{2\pi}{\lambda}s_{3}},\,\ldots,\,e^{\mathrm{j}\frac{2\pi}{\lambda}s_{N}}\end{bmatrix}^{\mathrm{T}},$ (4)
where $s_{n}\triangleq r_{n}-r_{1}$ denotes the path-length difference between the $n$th element and the reference (first) element, with $r_{n}$ being the distance from the UE to the $n$th active RIS element.
@@
$\mathbf{a}(\varphi,\phi)=\Big[$ $1,\,\ldots,\,e^{\mathrm{j}\tfrac{2\pi}{\lambda}\left(i(n)\sin\varphi\cos\phi+k(n)\sin\phi\right)},$ (5)
$\ldots,\,e^{\mathrm{j}\tfrac{2\pi}{\lambda}\left((N_{\mathrm{H}}{-}1)\Delta_{\mathrm{H}}\sin\varphi\cos\phi+(N_{\mathrm{V}}{-}1)\Delta_{\mathrm{V}}\sin\phi\right)}\Big]^{\mathrm{H}},$
where $i(n)$ and $k(n)$ are the horizontal and vertical coordinates (index positions) of the $n$th active RIS element, $\Delta_{H}$ and $\Delta_{V}$ are the inter-element spacings along the horizontal and vertical axes, and $N_{H}$ and $N_{V}$ denote the numbers of elements per row and column, respectively. We can express $i(n)$ and $k(n)$ in terms of $\Delta_{H}$ and $\Delta_{V}$ [12],
$i(n)=(n_{H}-1)\Delta_{H},\quad k(n)=(n_{V}-1)\Delta_{V},$
where,
$n_{H}=\text{mod}(n-1,N_{H})+1,\quad n_{V}=\left\lceil\frac{n}{N_{H}}\right\rceil.$
@@ III MLE-based Estimation for active RIS
We consider the received-signal model when the UE transmits a pilot symbol as
@@
where $\bm{y}\in\mathbb{C}^{L}$ is the received vector across $L$ pilot instances, $\bm{B}\in\mathbb{C}^{L\times N}$ is the active RIS configuration matrix over these instances, $P_{p}$ is the pilot power. We are sending the same pilot but for $L$ slots.
The deterministic mean is $\bm{\mu}(\bm{g})\;=\;\big(\bm{B}\,\mathbf{D}_{h}\bm{g}\big)\,\sqrt{P_{p}},$ and the effective noise is $\tilde{\bm{w}}\;=\;\bm{B}\,\mathbf{D}_{h}\bm{v}\;+\;\bm{w};\qquad\tilde{\bm{w}}\sim\mathcal{CN}\!\left(\bm{0},\,\bm{F}\right),$ with total noise covariance
@@
$f_{\bm{Y}}(\bm{y};\bm{g})=\frac{1}{\pi^{L}\det(\bm{F})}\,\exp\left(-\big(\bm{y}-\bm{\mu}(\bm{g})\big)^{\mathrm{H}}\bm{F}^{-1}\big(\bm{y}-\bm{\mu}(\bm{g})\big)\right)$ (8)
@@
$=\bm{y}^{\mathrm{H}}\bm{F}^{-1}\bm{y}-\bm{y}^{\mathrm{H}}\bm{F}^{-1}\bm{\mu}(\bm{g})-\bm{\mu}^{\mathrm{H}}(\bm{g})\bm{F}^{-1}\bm{y}+\bm{\mu}^{\mathrm{H}}(\bm{g})\bm{F}^{-1}\bm{\mu}(\bm{g})$
$=\bm{y}^{\mathrm{H}}\bm{F}^{-1}\bm{y}-2\,\Re\left\{\bm{y}^{\mathrm{H}}\bm{F}^{-1}\bm{\mu}(\bm{g})\right\}+\bm{\mu}^{\mathrm{H}}(\bm{g})\bm{F}^{-1}\bm{\mu}(\bm{g})$
@@ III-A ARIS Reflected Channel Phase $(\omega)$
The expression in (11) is minimized when the real part inside the objective is maximized. Therefore, the optimal phase $\hat{\omega}$ aligns the exponential term with the phase of the complex inner product.
@@
Now, by substituting the value of $\hat{\omega}$ into the equation (11) , we get
@@ III-B ARIS Reflected Channel Amplitude $(\beta)$
This is a quadratic function in terms of $\sqrt{\beta}$, for which the minimizer can be determined in closed form as
@@
III-C ARIS Reflected Channel Array Response Parameter ($\psi$)
By replacing this value $\hat{\beta}$ in equation (13), we get:
@@ IV Adaptive Beam and Amplification Factor
We employ an orthogonal angle-based codebook in place of the conventional DFT codebook. Let ${\Theta}=\{\bm{\theta}_{k}\in\mathbb{C}^{N}:\|\bm{\theta}_{k}\|_{\infty}=1,\;k=1,\ldots,K\}$ collect unit-modulus, phase-only beams generated from orthogonal angle pairs (far field: $(\varphi,\phi)$; near field: $(\varphi,\phi,r)$). By construction, $\bm{\theta}_{i}^{\mathrm{H}}\bm{\theta}_{j}\approx 0$ for $i\neq j$, so the search space contracts from an $N\times N$ DFT matrix to $R\ll N^{2}$ nearly orthogonal beams, each column representing a single active RIS configuration direction.
Initialization via wide-beam probing
Rather than selecting an arbitrary codebook column, we first form two wide beams that cover complementary sectors ( [13] section VI). Let $\bm{\theta}_{w,1}$ and $\bm{\theta}_{w,2}$ denote the two unit-modulus phase vectors. Each vector is configured to generate a wide beam covering a distinct angular region on either side of the active RIS, facilitating initial user discovery. The initial active RIS coefficients apply equal per-element gain with phases from the wide beams,
$\bm{\phi}_{1}\;=\;\sqrt{\tfrac{P_{\mathrm{RIS}}}{N}}\,\bm{\theta}_{w,1},\qquad\bm{\phi}_{2}\;=\;\sqrt{\tfrac{P_{\mathrm{RIS}}}{N}}\,\bm{\theta}_{w,2},$ (16)
Here, $\bm{\phi}_{1}$ and $\bm{\phi}_{2}$ are the initial active RIS configurations. The matrix $\mathbf{B}$ initially contains only these two configurations. Each time a pilot is transmitted, the newly selected active RIS configuration is appended to $\mathbf{B}$ as an additional row. These pilots reliably illuminate the UE in both near- and far-field regimes and significantly reduce the required pilot count in the first step.
Active-RIS amplitude shaping and phase alignment:
Given $\widehat{\bm{g}}$, we design the per-element phases for coherent combining across the cascaded links:
@@
which cancels the aggregate phase $h_{n}\widehat{g}_{n}$ at each element. The per-element amplification vector $\bm{p}$ is then computed in closed form under the active RIS power constraint using [14, Sec. V, Cor. 3] given by,
@@
where,
$\alpha_{n}$ $=|\widehat{g}_{n}|\,|h_{n}|,\qquad\beta_{n}=|h_{n}|^{2},\qquad\gamma_{n}=\frac{|\widehat{g}_{n}|^{2}\,\dfrac{p_{d}}{\sigma_{v}^{2}}+1}{\dfrac{P_{\mathrm{ris}}}{\sigma^{2}}},$ (19)
$C$ $=\left(\sum_{n=1}^{N}\frac{|\alpha_{n}|^{2}\,\gamma_{n}}{(\beta_{n}+\gamma_{n})^{2}}\right)^{-1/2}.$
For the initial pilots, the channel is unknown; therefore, this amplification factor cannot be applied. After processing initial pilots and obtaining a channel estimate, we update the amplification factor in the next active RIS configuration. The resulting active RIS configuration, $\bm{\phi^{\star}}$ is then used for pilot transmission, which is
$\bm{\phi}^{\star}\;=\;\bm{p}^{(\mathrm{est})}\odot\bar{\bm{\theta}}^{\mathrm{opt}},$ (20)
@@
1:Input: $\mathbf{D}_{h}=\mathrm{diag}(\mathbf{h})$, codebook $\bm{\Theta}$ (phase-only), initial configurations $\{\bm{\theta}_{w,1},\bm{\theta}_{w,2}\}$, number of slots $L$, powers $P_{p}$, $P_{d}$, $P_{\text{RIS}}$
2:Initialization: Set $\mathbf{D}_{h,\angle}\leftarrow\mathrm{diag}(\mathbf{h}/|\mathbf{h}|)$ $\bm{\phi}_{1}\leftarrow\sqrt{P_{\mathrm{RIS}}/N}\,\bm{\theta}_{w,1}$,  $\bm{\phi}_{2}\leftarrow\sqrt{P_{\mathrm{RIS}}/N}\,\bm{\theta}_{w,2}$

### v2

@@ I Introduction
Next-generation (6G) wireless networks require mechanisms that reconfigure the propagation environment to improve end-to-end performance. Reconfigurable intelligent surfaces (RISs) address this need by enhancing wireless channels via programmable metasurfaces [15]. A passive RIS uses many low-cost scatterers to reflect incident signals with adjustable phase shifts, shaping the channel without extra transmit power [6]. However, passive RIS suffers from a double-fading effect [9], motivating architectures that can compensate for the severe cascaded path loss. To overcome this limitation, active RIS (ARIS) architectures integrate active components that amplify rather than merely reflect the impinging wave.
Even with these advantages, reliably acquiring channel state information (CSI) in RIS-aided systems remains challenging. Conventional schemes require the Base Station/User to transmit pilots while the RIS cycles through many configurations, so the pilot typically scales with the number of elements, which becomes prohibitive for large surfaces and under mobility [13]. Mobility-aware designs based on statistical temporal correlation can track small-scale fading in localized regions, but they assume stationary statistics and still require pilot lengths on the order of the RIS size, making them unsuitable for non-stationary motion or large coverage areas [14]. Subarray-based reduced subspace methods can shrink the effective training dimension, yet either sacrifice beamforming gain or still incur substantial overhead for physically large RISs [4]. Parametric approaches exploit the low-rank of mmWave channels to reduce cascaded-channel estimation into angle and gain estimation problems using matrix factorization, compressive sensing, or atomic norm minimization but are mainly developed for line-of-sight (LoS) dominant far-field scenarios [8]. Recent near-field works for XL-MIMO and RIS for THz systems introduce distance aware parametric models and polar-domain sparsity. However, they are still in an early stage and often treat only the RIS path [3].
Building on these observations, Haghshenas et al. proposed in [7] a parametric MLE framework that directly targets most of these issues for passive RISs. Their method jointly estimates the direct BS–user channel and the LoS user–RIS channel, recovering angular parameters in the far field and, in the near field, the user–RIS distance, so that both near and far-field regimes are handled within a single parametric model. To curb pilot overhead, they combine this MLE with an adaptive RIS configuration strategy, and a wide-beam initialization further ensures useful observations even at low SNR and under mobility. The framework is, however, developed under the standard ideal passive RIS assumptions: each RIS element is modeled as a lossless phase shifter with perfect control, and the signal model includes only the usual receiver thermal noise, which is appropriate for a purely passive surface. Even with such improved channel acquisition, a passive RIS link remains fundamentally limited by the multiplicative fading effect. To overcome this limitation, one needs active surfaces that compensate for it by amplifying the reflected signal.
Very recently, researchers have shifted their focus to active RIS-aided CSI. Authors in [11] applied MLE to an active RIS-aided LoS channel for user localization, showing that parametric modeling can achieve high accuracy with active surfaces. The authors in [2] analyzed hardware impairment effects by deriving the Cramér–Rao bound for active RIS systems and proposing a joint least-squares channel estimator. In both these cases, the primary focus was the mean-squared error (MSE) rather than pilot overhead. In a complementary direction, Lee et al. [10] studied a semi-passive RIS with a few active sensing elements and reconstructed the full RIS–user channel via a correlation-aware linear combination of partial CSI, so that the training overhead no longer scales with the total number of RIS elements. However, neither does it exploit the per-element amplification nor the gain and phase calibration constraints.
Motivated by these gaps, we propose a parametric channel estimation methodology tailored to active RIS architectures. Specifically, we formulate a maximum likelihood estimator that yields closed-form estimates for all relevant channel parameters under both near and far-field conditions. The estimator jointly recovers the full complex gain (amplitude and phase) of the channel and explicitly obtains the user’s geometric parameters: angles of arrival and departure for far-field users and angles plus distance for near-field users within a unified framework. We further integrate an adaptive training and beamforming strategy that iteratively refines the RIS configuration based on initial estimates, concentrating pilot power in the most informative directions. This active RIS design aligns both amplification and phase shifts with the parametric estimation objective, thereby improving accuracy without increasing pilot length. To the best of our knowledge, this is the first comprehensive parametric channel estimator developed specifically for active RIS hardware, accounting for amplified noise and operating seamlessly across near and far-field regimes.
@@ II System and Channel Model
We consider an uplink system comprising a single-antenna base station (BS) and a single-antenna user equipment (UE) assisted by an active reconfigurable intelligent surface (ARIS). No direct BS-to-UE link is assumed. The active RIS comprises $N$ elements with independently tunable amplitude and phase. We model the BS-to-ARIS channel, $\bm{h}\in\mathbb{C}^{N}$, as a deterministic LoS link. Given that the BS and active RIS are fixed in location, this channel is considered quasi-stationary and non-fading over the signaling interval. The ARIS-to-UE channel is $\bm{g}\triangleq[g_{1},\ldots,g_{N}]^{\mathrm{T}}\in\mathbb{C}^{N}$, which is time varying due to user mobility. To maintain a high signal-to-noise ratio (SNR), the BS periodically estimates the channel and updates the active RIS coefficients (phase shifts and gains) accordingly.
Let the UE transmit the symbol $x\sim\mathcal{CN}(0,P_{d})$. The received signal at the BS is:
@@
where $\bm{\Phi}\triangleq[\Phi_{1},\ldots,\Phi_{N}]^{\mathrm{T}}$ , it collects the complex active RIS coefficients. Here $\mathbf{D}_{h}=\operatorname{diag}(h_{1},\ldots,h_{N})$, and $\Phi_{n}=p_{n}e^{\mathrm{j}\theta_{n}}$ is the active RIS configuration, where $\theta_{n}$ is the phase shift and $p_{n}$ is the amplification factor of the $n$th element, respectively.
@@
The spectral efficiency (SE) can be expressed as:
@@
The LoS channel vector from UE-to-ARIS is modeled as:
$\bm{g}\;=\;\sqrt{\beta}\,e^{\mathrm{j}\omega}\,\bm{a}(\psi),$ (3)
where $\beta\geq 0$ and $\omega\in[0,2\pi)$ denote, respectively, the channel gain and the phase referenced to the first active RIS element, and $\bm{{a}(\psi)}$ is the array-response vector parameterized by $\bm{\psi}$. This model applies to both near and far-field propagation: in the far field, $\bm{\psi}=(\varphi,\phi)$ comprises the azimuth $\varphi$ and elevation $\phi$; in the near field, $\bm{\psi}=(\varphi,\phi,r)$ additionally includes the distance $r$ from the UE to the first active RIS element (Reference element).
For the near field, the array response vector can be written as:
$\mathbf{a}(\varphi,\phi,r)=\begin{bmatrix}1,\,e^{-\mathrm{j}\frac{2\pi}{\lambda}s_{2}},\,e^{-\mathrm{j}\frac{2\pi}{\lambda}s_{3}},\,\ldots,\,e^{-\mathrm{j}\frac{2\pi}{\lambda}s_{N}}\end{bmatrix}^{\mathrm{T}},$ (4)
where $s_{n}\triangleq r_{n}-r_{1}$ denotes the path-length difference between the $n$th element and the reference element, with $r_{n}$ being the distance from the UE to the $n$th active RIS element.
@@
$a(\varphi,\phi)=\bigg[$ $1,\ldots,e^{-j\frac{2\pi}{\lambda}\left[i(n)\Delta_{\mathrm{H}}\sin(\varphi)\cos(\phi)+j(n)\Delta_{\mathrm{V}}\sin(\phi)\right]},$ (5)
$\ldots,e^{-j\frac{2\pi}{\lambda}\left[(N_{\mathrm{H}}-1)\Delta_{\mathrm{H}}\sin(\varphi)\cos(\phi)+(N_{\mathrm{V}}-1)\Delta_{\mathrm{V}}\sin(\phi)\right]}\bigg]^{\mathrm{T}},$
where $i(n)$ and $j(n)$ are the horizontal and vertical coordinates of the $n$th active RIS element, $\Delta_{H}$ and $\Delta_{V}$ are the inter-element spacings along the horizontal and vertical axes, and $N_{H}$ and $N_{V}$ denote the numbers of elements per row and column, respectively. We can express $i(n)$ and $j(n)$ in terms of $\Delta_{H}$ and $\Delta_{V}$ [1], $i(n)=\operatorname{mod}(n-1,N_{\mathrm{H}})\quad\text{and}\quad j(n)=\left\lfloor\frac{n-1}{N_{\mathrm{H}}}\right\rfloor$.
@@ III MLE-based Estimation for active RIS
We consider the received signal model when the UE transmits a pilot symbol as:
@@
where $\bm{y}\in\mathbb{C}^{L}$ is the received vector across $L$ instances of the pilot, $\bm{B}\in\mathbb{C}^{L\times N}$ is the active RIS configuration matrix over these instances, $P_{p}$ is the pilot power. We are sending the same pilot but for $L$ slots.
The deterministic mean is $\bm{\mu}(\bm{g})\;=\;\big(\bm{B}\,\mathbf{D}_{h}\bm{g}\big)\,\sqrt{P_{p}},$ and the effective noise is $\tilde{\bm{w}}\;=\;\bm{B}\,\mathbf{D}_{h}\bm{v}\;+\;\bm{w};\qquad\tilde{\bm{w}}\sim\mathcal{CN}\!\left(\bm{0},\,\bm{F}\right),$ with total noise covariance,
@@
$f_{\bm{Y}}(\bm{y};\bm{g})=\frac{1}{\pi^{L}\det(\bm{F})}\,\exp\left[-\big(\bm{y}-\bm{\mu}(\bm{g})\big)^{\mathrm{H}}\bm{F}^{-1}\big(\bm{y}-\bm{\mu}(\bm{g})\big)\right].$ (8)
@@ III-A ARIS Reflected Channel Phase $(\omega)$
The expression in (11) is minimized when the real part inside the objective is maximized. Therefore, the optimal phase $\hat{\omega}$ aligns the exponential term with the phase of the complex inner product as:
@@
Now, by substituting the value of $\hat{\omega}$ into the equation (11) , we get:
@@ III-B ARIS Reflected Channel Amplitude $(\beta)$
This is a quadratic function in terms of $\sqrt{\beta}$, for which the minimizer can be determined in closed form as:
@@
III-C ARIS Reflected Channel Array Response Parameter ($\bm{\psi}$)
By replacing this value $\hat{\beta}$ in equation (13), we get
@@ IV Adaptive Beam and Amplification Factor
Codebook & Search Space: Let $\Theta=\{\bm{\theta}_{k}\in\mathbb{C}^{N}:\ \|\bm{\theta}_{k}\|_{\infty}=1,\ k=1,\ldots,K\}$ collect phase-only beams generated from orthogonal angle pairs (Same codebook can be used for near field too, according to Weyl identity [12]). By construction, $\bm{\theta}_{i}^{H}\bm{\theta}_{j}\approx 0$ for $i\neq j$, so the search contracts from an $N\times N$ DFT matrix to $R\ll N^{2}$ nearly orthogonal beams, each column representing a single RIS configuration direction.
Initialization via Wide-Beam Probing: Following [7, Sec. VI], two complementary wide beams with unit-modulus phase vectors $\bm{\theta}_{w,1}$ and $\bm{\theta}_{w,2}$ is designed. The active RIS is configured with equal per-element gain and the wide-beam phases as:
$\bm{\Phi}_{1}=\sqrt{\frac{P_{\mathrm{RIS}}}{N}}\ \bm{\theta}_{w,1},\qquad\bm{\Phi}_{2}=\sqrt{\frac{P_{\mathrm{RIS}}}{N}}\ \bm{\theta}_{w,2}.$ (16)
The configuration matrix B is initialized with these two rows; each new pilot appends its active RIS configuration as an additional row. These pilots illuminate the UE in both near and far-field regimes and reduce the initial pilot count. $P_{\mathrm{RIS}}$ represents the total allowable active RIS power.
Active-RIS Amplitude Shaping & Phase Alignment With $\bm{\hat{g}}$, per-element phases for coherent combining across the cascaded links is taken by
@@
which cancels the aggregate phase $h_{n}\widehat{g}_{n}$ at each element. Under the active-RIS power constraint using [5, Sec. V, Cor. 3], the per-element amplification can be computed as:
@@
where, $\alpha_{n}=|\widehat{g}_{n}|\,|h_{n}|,$ $\beta_{n}=|h_{n}|^{2},$ $\gamma_{n}=\tfrac{|\widehat{g}_{n}|^{2}\,\tfrac{p_{d}}{\sigma_{v}^{2}}+1}{\tfrac{P_{\mathrm{ris}}}{\sigma^{2}}},$ $C=\Big(\sum_{n=1}^{N}\frac{|\alpha_{n}|^{2}\,\gamma_{n}}{(\beta_{n}+\gamma_{n})^{2}}\Big)^{-1/2}.$
For the initial pilots, the channel is unknown, so this amplification is not applied. After estimating the channel from the initial pilots, the amplification factors are updated and transmitted through the next pilot using
$\bm{\Phi}^{\star}=\bm{p}^{(\mathrm{est})}\odot\bar{\bm{\theta}}_{\mathrm{opt}},$ (20)
@@
Adaptive Pilot Selection The estimate can be refined by choosing the next phase-only codeword that best aligns with the current active vector
and set
$\bm{\Phi}_{\mathrm{new}}=\bm{p}^{(\mathrm{est})}\odot\bm{\theta}_{\mathrm{next}}.$ (22)
We append $\bm{\Phi}_{\text{new}}$ as a new row to the configuration matrix $\mathbf{B}$, collect the corresponding measurement, and update $\widehat{\bm{g}}$ via the MLE step. This procedure is repeated greedily until the pilot budget is exhausted. Each selected codeword is removed from $\Theta$.
Complexity & Applicability Each refinement evaluates the remaining phase-only codewords $\Theta_{\mathrm{rem}}$ using $|\Theta_{\mathrm{rem}}|$ length-NN inner products, i.e., $\mathcal{O}\!\left(N|\Theta_{\mathrm{rem}}|\right)$ per iteration, for a total training complexity $\mathcal{O}(LNR)$ with $L$ pilots and codebook size $R$ ($R\ll N^{2}$), greatly reducing cost versus DFT exhaustive search.
@@
1:codebook $\Theta$ (phase-only), initial wide beams $\{\bm{\theta}_{w,1},\bm{\theta}_{w,2}\}$, slots $L$, powers $(P_{p},P_{d},P_{\rm RIS})$, $\mathbf{D}_{h}=\mathrm{diag}(\mathbf{h})$